OpenStudy (anonymous):

write an equation of the line containing the given point and parellel to the given line (5,9);x+8y=7

OpenStudy (aroub):

y=mx+b ,m=-1/8 , (5,9) 9=-1/8(5)+b 9=-5/8 +b 5/8+72/8=b 77/8 =b y=-1/8x+77/8

OpenStudy (anonymous):

so it y= 1/8x+77/8 right

OpenStudy (aroub):

-1/8x

OpenStudy (aroub):

yes..

OpenStudy (anonymous):

thank you i though there was no - in there sorry but thanks

OpenStudy (anonymous):

i got the same flower

OpenStudy (aroub):

your welcome.. / i used to have the same flower too then i changed the pic... lol